1. An underwater swimmer looks up toward the surface of the water on a line of sight that makes an angle of 25˚ with a normal to the surface of the water. What is the angle of incidence in air for the light rays that enter the swimmer’s eye?



2. In each of the following questions, the second medium is air.

a) What is the critical angle if the index of refraction for a medium is 1.68?

b) What is the index of refraction of a medium if the critical angle is 40˚?


1
Expert's answer
2022-05-30T15:17:08-0400

1.

"\alpha=36°,"

2.

"\alpha_1=\arcsin \frac 1{ n_1}=37°,"

"n_2=\frac 1{\sin\alpha_2}=1.56."